### v1

# RobustFormer: Noise-Robust Pre-training for Images and Videos

Abstract
While deep learning models are powerful tools that revolutionized many areas, they are also vulnerable to noise as they rely heavily on learning patterns and features from the exact details of the clean data. Transformers, which have become the backbone of modern vision models, are no exception. Current Discrete Wavelet Transforms (DWT) based methods do not benefit from masked autoencoder (MAE) pre-training since the inverse DWT (iDWT) introduced in these approaches is computationally inefficient and lacks compatibility with video inputs in transformer architectures.
In this work, we present RobustFormer, a method that overcomes these limitations by enabling noise-robust pre-training for both images and videos; improving the efficiency of DWT-based methods by removing the need for computationally iDWT steps and simplifying the attention mechanism. To our knowledge, the proposed method is the first DWT-based method compatible with video inputs and masked pre-training. Our experiments show that MAE-based pre-training allows us to bypass the iDWT step, greatly reducing computation. Through extensive tests on benchmark datasets, RobustFormer achieves state-of-the-art results for both image and video tasks.
Keywords Noise Robustnesss, Discrete Wavelet Transform, Masked Pre-training
1 Introduction
Addressing robustness in deep learning models for video data is crucial for practical applications, including surveillance, autonomous driving, and multimedia content analysis [34]. Just as images can suffer from distortions such as blurring, noise, and artifacts, video data experiences similar issues, compounded by the additional variability introduced over time [38]. Deep learning models, especially those trained on clean, high-quality data, are highly sensitive to these disruptions because they rely heavily on precise patterns within the data to make accurate predictions [48]. Even minor noise or inconsistencies in a single frame can cause the model to misinterpret important features, as these models have not been trained to ignore unexpected distortions [42]. With video data, the challenge becomes even greater due to the temporal nature of the information, where each frame builds on previous ones to create a coherent sequence [54]. For instance, temporal inconsistencies such as random noise across frames, motion blur, flickering shadows, and heat-induced turbulence can lead to abrupt changes between frames, confusing the model as it attempts to track and interpret motion or recognize actions [15, 54]. In tasks like autonomous driving, these inconsistencies are highly prevalent and pose significant risks, as the model may fail to accurately detect objects, assess distances, or predict movement patterns [40].
A primary reason for these issues lies in the pixel-based design of most deep-learning models. Although RGB pixel representations are the standard in vision applications because they capture detailed color and spatial information, they are computationally demanding and highly sensitive to noise and domain shifts. This sensitivity arises since RGB data represents exact pixel-level details, so even slight distortions can disrupt deep-learning models that depend on these patterns for accuracy. Additionally, RGB representations process both essential and irrelevant features equally, limiting their efficiency. Inspired by image compression techniques like JPEG, frequency domain transformations such as Discrete Cosine Transform (DCT) provide a solution by isolating low-frequency components (representing significant information) from high-frequency ones (less relevant details) [32, 39]. This separation reduces memory demands and increases robustness to noise without compromising performance [11, 16, 32].
While pure frequency domain transformations offer these benefits, they also lead to a loss of spatial or temporal locality due to the change in basis, a consequence of the Heisenberg uncertainty principle. In contrast, Discrete Wavelet Transform (DWT) provides a more balanced approach, decomposing the signal into both high- and low-frequency components across multiple scales. This multi-resolution representation allows DWT to retain spatial and temporal information, preserving important features and contextual details while still isolating high-frequency information. Image-based methods leveraging DWT have shown promise in enhancing robustness against spatial noise by allowing models to focus on frequency-based features that are less affected by disruptions [30]. Applying these principles to video data can address both temporal and spatial corruptions, as DWT can help models focus on low-frequency information, improving robustness to noise and other distortions across frames.
In this paper, we employ a Video Masked Autoencoding approach [49, 50], additionally, incorporating DWT for pre-training and fine-tuning on the uncorrupted datasets, followed by evaluation on their corrupted counterparts. We use UCF101 and HMDB51 datasets for training and evaluation, where the generated corruptions are of 6 types(including 5 spatial and 1 temporal corruptions in 5 severity levels) reflecting real world scenarios [57].
In summary, our contributions are as follows:
1.
We present a novel Discrete Wavelet Transform (DWT) based masked autoencoder architecture that is robust to spatial and temporal corruptions in both video and image data. To our knowledge this is the first work to implement DWT in Masked autoencoder setting.
2.
We perform comprehensive evaluation on several real-world noise types with varying severity levels in large-scale benchmark datasets.
3.
We demonstrate that our method performs equally well and in many cases better than commonly used IDWT variant which requires comparatively more compute resource and makes architectures overly complicated.
2 Related Works
2.1 Masked Video Training
BERT introduced the concept of Masked Language Modeling (MLM), a novel pretraining objective for natural language understanding, which led to significant advances in NLP [26]. Following BERT’s success, various NLP models, including RoBERTa [35], ALBERT [28], and ELECTRA [6], extended this masked training approach. Inspired by BERT’s masked token prediction, the concept of Masked Image Modeling (MIM) emerged for computer vision tasks. iGPT [5] first adapted the Transformer model for image data by treating images as pixel sequences and using a similar masked prediction task. BEiT [1] extended this concept by proposing a discrete variational autoencoder (dVAE) to tokenize image patches, predicting masked tokens in a BERT-style pretraining. These approaches demonstrated that MIM could effectively learn transferable representations for various downstream tasks. Masked Autoencoders (MAE) [19] refined MIM by reconstructing masked regions in images, achieving state-of-the-art results on multiple benchmarks using Vision Transformers (ViTs) and large-scale unlabeled data.
Building on masked training in language and image domains, Masked Video Modeling (MVM) has emerged to capture both temporal and spatial representations from video data. VideoMAE [49] extends MAE to videos by applying random masking across both spatial and temporal dimensions, enabling the model to learn robust spatio-temporal representations, as shown in Figure 2 (a). VideoMAE achieves state-of-the-art performance on numerous video classification tasks, highlighting the effectiveness of masked training for video data. Similarly, BEVT (BEiT for Video) [51] extends BEiT’s tokenization and masked prediction strategy to video patches, achieving strong results in action recognition tasks.
2.2 Discrete Wavelet transform(DWT)
Wavelets are essential in time-frequency analysis and signal processing tasks, such as anti-aliasing and detail restoration, through Discrete Wavelet Transform (DWT) and its inverse (IDWT) [36]. Early studies combined wavelets with shallow neural networks for function approximation and classification, optimizing wavelet parameters within the network [47]. Recently, deeper networks have adopted wavelets for image classification, although these integrations can be computationally intensive [7]. The Multi-level Wavelet CNN (MWCNN) [33] applies Wavelet Packet Transform (WPT) for image restoration, handling both low- and high-frequency components. Similarly, the Convolutional-Wavelet Neural Network (CWNN) [10] utilizes dual-tree complex wavelet transform (DT-CWT) to reduce noise in SAR images while preserving crucial features, though within a simplified two-layer structure. Wavelet Pooling [53] employs a two-level DWT for pooling and combines DWT/IDWT for backpropagation, deviating from conventional gradient methods.
A key advantage of wavelets is their ability to provide sparse representations, which enables more efficient processing and faster model training [53, 31, 54]. DWT-based training reduces redundancy, making it well-suited for real-time video classification tasks where small shifts in data can significantly affect model output. Unlike other approaches, our method avoids reconstruction after the wavelet transform, thus saving considerable computational resources, especially when handling high-resolution video data. This streamlined process enhances efficiency while maintaining robust performance in noisy environments.
2.3 Noise Robustness
Robustness in video classification has been an ongoing research focus, initially centered on adapting image robustness techniques to video data. For instance, [22] introduced benchmarks for evaluating image classifiers’ robustness against corruptions, a framework later adapted to videos by [24] to assess model performance under various perturbations. Addressing the challenge of temporal consistency in video data, models such as the two-stream network [44] and 3D convolutions [3] were developed, with data augmentation methods like temporal jittering and frame dropping [43] further enhancing robustness.
Since DWT has been widely applied in noisy data reconstruction, achieving notable gains in both noise removal and complex tasks like deblurring [53, 31], our approach builds on the methodologies of WaveFormer [55] and WaveCNets [30] by incorporating DWT. As illustrated in Figure 2, (b) shows WaveFormer’s framework, which computes attention using only the low-frequency components of the query and key matrices. Unlike WaveFormer, however, we omit DWT and inverse DWT (IDWT) for the value matrix, passing it directly. Similar to WaveCNets, we discard high-frequency components during the initial downsampling layer but extend this technique to 3D-DWT to handle temporal decomposition.
3 Methods
Compared to the classical VideoMAE [48], our proposed method enhances robustness by employing 3D-DWT for spatio-temporal analysis; in contrast to WaveFormer [54], our approach eliminates the computationally expensive iDWT step, significantly improving efficiency, as illustrated in Figure 2.
Our implementation incorporates a two-step wavelet transform. First, we utilize DWT’s efficient downsampling capability to generate embeddings for video patches. By leveraging 3D-DWT, we can process video data more effectively than stacking coefficients from 2D-DWT. This approach enables simultaneous handling of both temporal and spatial noise, ensuring robust feature extraction. The forward and backward propagation processes for this step are detailed below.
3.1 Wavelet Embedding
Let be the original downsampled and clipped video sequence of sequence length , height , width and channels . We start with the generation of transformation matrices based on the selected wavelet filters. This step is critical as it defines the low-pass and high-pass filters that are applied in the transformation process. Wavelet Filter Coefficients: Define the wavelet by its low-pass (L) and high-pass (H) filter coefficients.
Now we construct the transformation matrices for each dimension (depth, height, and width). For simplicity, only the depth dimension’s matrix construction is shown here. We can replicate this for the height and width dimensions:
where, k {H, W, D} represent each of the individual dimension. The forward pass thus involves applying these matrices to decompose the input data into its respective frequency sub-bands. The input tensor is thus decomposed into eight sub-bands using the transformation matrices for depth , height , and width , as shown in Figure 3.
where, the matrix construction leads to generating matrix for dimension corresponding to . The resulting matrix will have dimensions ,where is the ceiling function that handles cases where is odd. for filter and . The ConstructMatrix function works as follows:
where, where indexes the filter coefficients .
3.2 Attention Computation
In the wavelet embedding step we average, concat or omit the high frequency components in an effort to avoid noise and smooth out the input representation. However since we perform additional latent transformation(3D Convolution with tube masking) which might escalate some of the remaining noises during attention calculation. We want to minimize the redundant details as much as possible during our overall training. Prior works[54] have proved that noise before the attention mechanism alters the correlation scores between features by increasing the values for unrelated pairs (due to the random alignment by noise) and decreasing them for related pairs (as noise dilutes the true alignment), thereby distorting the focus of the attention mechanism. This in our case leads to highly deviated classification accuracy. We thus follow steps from [54] to smooth out the latent representation before proceeding to attention calculation by completely omitting the involved high frequency components.
Suppose denotes the output of the tube-masked 3D convolution, representing the deep encoded feature representations. Here, each is a tensor in , where is the tubelet size, is the embedding size, and is the patch size. The elements of the attention mechanism, , , and , are first calculated by passing through three different linear layers.
We then perform 1D-DWT for each Q, K and V by obtaining the low and high pass filters corresponding to a specific wavelet(for eg. haar) similar to [54] as follows,
We calculate these segregated components similar to 2 and omit and . Note we also omit dwt computation for , so that the attention computation can attend to some useful high level components from (that are omitted in and in the input, as shown in Figure 3. With this we can avoid performing additional IDWT step as in [54] which makes the architecture complicated and computationally intensive.
We finally perform our attention score computation as follows:
The backward function is defined for each dwt computation similar to 3D dwt computation.
With this we create various configurations of Robustformer that involves various operations over DWT blocks. RF-A indicates Robustformer with dwt averaging, RF-AA refers to averaging as well as dwt attention as mentioned in 5. Similarly we define RF-O, RF-OA, RF-C and RF-CA for omit, omit using dwt attention, concat and concat using dwt attention.
Models | IN-C (mCE) |
CNNs | |
ResNet-50 [21] | 76.7 |
ResNeXt50-32x4d [56] | 64.7 |
VGG-11 [45] | 93.5 |
VGG-19 [45] | 88.9 |
VGG-19 + BN [45] | 81.6 |
ANT [41] | 63.0 |
EWS [17] | 63.0 |
WResNet-18* [31] | 75.6 |
WResNet-18* [31] | 65.8 |
ViTs | |
PVT-Large [52] | 59.8 |
BiT-mr101x3 [27] | 58.3 |
MAE-ViT-B [49] | 58.8 |
Robust Formers (Ours) | |
RobustFormer-A | 55.3 |
RobustFormer-AA | 55.7 |
RobustFormer-O | 55.3 |
RobustFormer-OA | 55.9 |
RobustFormer-C | 57.7 |
RobustFormer-CA | 58.3 |
Model | Params | Flops | IN-P(mFP) |
MAE-ViT-B [49] | 111.7M | 36.7G | 13.4 |
RobustFormer-A | 111.2M | 36.6G | 12.7 |
RobustFormer-AA | 111.2M | 36.6G | 12.7 |
RobustFormer-O | 111.6M | 37.5G | 12.2 |
RobustFormer-OA | 111.6M | 37.5G | 12.4 |
RobustFormer-C | 111.2M | 38.0G | 12.9 |
RobustFormer-CA | 111.2M | 38.0G | 12.9 |
Network | Noise | Blur | Weather | Digital | |||||||||||
Gauss. | Shot | Impulse | Defocus | Glass | Motion | Zoom | Snow | Frost | Fog | Bright | Contrast | Elastic | Pixel | JPEG | |
AlexNet | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 97 | 98 | 96 | 109 | 134 |
SqueezeNet | 107 | 106 | 105 | 103 | 101 | 100 | 100 | 101 | 103 | 97 | 98 | 98 | 100 | 98 | 106 |
VGG-11 | 97 | 97 | 100 | 92 | 93 | 91 | 94 | 94 | 94 | 89 | 84 | 75 | 84 | 78 | 87 |
VGG-19 | 92 | 89 | 91 | 89 | 89 | 89 | 90 | 89 | 89 | 86 | 78 | 64 | 80 | 84 | 85 |
VGG-19+BN | 82 | 83 | 88 | 84 | 84 | 83 | 89 | 82 | 87 | 85 | 68 | 59 | 87 | 74 | 85 |
ResNet-18 | 87 | 88 | 91 | 84 | 91 | 87 | 88 | 91 | 88 | 78 | 69 | 69 | 84 | 88 | 80 |
ResNet-50 | 80 | 82 | 83 | 75 | 89 | 78 | 80 | 78 | 78 | 75 | 66 | 57 | 71 | 85 | 87 |
WResNet-18* | 80.2 | 80.5 | 80.5 | 79.7 | 89.8 | 83.6 | 84.8 | 84.9 | 80.8 | 73.9 | 66.3 | 75.1 | 88.2 | 75.1 | 88.6 |
WResNet-101* | 64.3 | 65.9 | 65.0 | 63.6 | 80.4 | 68.2 | 71.8 | 71.6 | 67.5 | 60.2 | 50.1 | 62.9 | 74.6 | 52.3 | 67.9 |
AugMix | 67 | 66 | 68 | 64 | 79 | 59 | 64 | 69 | 68 | 65 | 54 | 57 | 74 | 60 | 66 |
MAE-ViT | 56.7 | 57.3 | 56.1 | 61.9 | 75.4 | 58.1 | 68.6 | 57.7 | 63.4 | 53.7 | 45.4 | 45.0 | 68.9 | 55.1 | 58.7 |
RF-O | 51.9 | 52.3 | 52.0 | 58.0 | 71.7 | 56.0 | 65.2 | 57.1 | 58.9 | 54.4 | 44.2 | 45.4 | 63.4 | 45.6 | 56.5 |
RF-OA | 51.6 | 51.8 | 51.8 | 59.2 | 72.2 | 55.9 | 64.8 | 56.9 | 60.2 | 55.8 | 44.4 | 45.3 | 64.2 | 47.9 | 56.4 |
RF-A | 52.9 | 53.5 | 52.4 | 58.7 | 68.7 | 55.9 | 64.4 | 56.6 | 58.6 | 55.3 | 44.5 | 43.7 | 61.6 | 47.8 | 55.2 |
RF-AA | 52.8 | 53.9 | 52.8 | 58.9 | 69.1 | 57.4 | 64.1 | 56.8 | 59.5 | 55.5 | 44.9 | 44.8 | 61.9 | 48.0 | 55.9 |
RF-C | 53.6 | 54.6 | 52.1 | 62.9 | 76.6 | 57.4 | 69.2 | 55.9 | 61.9 | 52.8 | 44.2 | 45.4 | 69.1 | 52.2 | 57.3 |
RF-CA | 54.6 | 55.6 | 53.4 | 62.8 | 75.9 | 57.2 | 69.2 | 57.4 | 62.6 | 53.3 | 44.4 | 47.1 | 69.4 | 53.5 | 57.5 |
Network | Params | Flops | Noise | Blur | Temporal | Digital | Camera | Mean | ||||||
R3D[18] | 32.5M | 55.1G | .71 | .61 | .78 | .70 | .98 | .97 | .91 | .88 | .89 | .85 | .85 | .80 |
I3D[4] | 28.0M | 75.1G | .72 | .61 | .80 | .72 | .97 | .96 | .91 | .87 | .89 | .85 | .86 | .80 |
SF[14] | 34.6M | 66.6G | .73 | .64 | .81 | .73 | .96 | .94 | .91 | .87 | .88 | .84 | .85 | .80 |
X3D[13] | 3.8M | 5.15G | .71 | .62 | .81 | .75 | .96 | .94 | .90 | .86 | .85 | .84 | .85 | .80 |
TF[2] | 122M | 196G | .87 | .84 | .91 | .87 | .98 | .96 | .94 | .94 | .95 | .93 | .91 | .88 |
MViT[12] | 36.6M | 70.7G | .93 | .91 | .86 | .82 | .96 | .95 | .94 | .93 | .92 | .92 | .93 | .91 |
VideoMAE[49] | 94.8M | 167.7G | .95 | .92 | .89 | .79 | .97 | .95 | .97 | .95 | .84 | .69 | .92 | .86 |
RF-A | 93.2M | 162.8G | .94 | .90 | .90 | .82 | .97 | .95 | .97 | .94 | .85 | .69 | .93 | .86 |
RF-AA | 93.2M | 162.8G | .94 | .91 | .90 | .80 | .96 | .95 | .97 | .95 | .85 | .70 | .92 | .86 |
RF-O | 93.2M | 160.4G | .94 | .91 | .90 | .81 | .97 | .97 | .96 | .96 | .96 | .68 | .95 | .87 |
RF-OA | 93.2M | 160.4G | .95 | .91 | .90 | .80 | .97 | .96 | .97 | .94. | .86 | .70 | .93 | .86 |
Perturbation | R3D | I3D | SF | X3D | TF | MViT | VideoMAE | RF-A | RF-AA | RF-OA |
Defocus Blur | .77 | .75 | .80 | .85 | .80 | .83 | .87 | .88 | .87 | .88 |
Motion Blur | .63 | .60 | .64 | .66 | .75 | .82 | .85 | .89 | .89 | .88 |
Zoom Blur | .71 | .74 | .80 | .85 | .89 | .90 | .65 | .65 | .65 | .65 |
Gaussian | .47 | .46 | .36 | .49 | .75 | .87 | .88 | .85 | .85 | .87 |
Shot | .78 | .79 | .82 | .79 | .94 | .95 | .96 | .96 | .98 | .96 |
Impulse | .44 | .42 | .34 | .46 | .76 | .87 | .87 | .83 | .84 | .86 |
Speckle | .75 | .75 | .70 | .75 | .91 | .95 | .95 | .95 | .95 | .96 |
Compression | .93 | .90 | .92 | .89 | .94 | .94 | .95 | .95 | .95 | .96 |
Static Rotate | .67 | .65 | .70 | .71 | .82 | .87 | .85 | .86 | .86 | .87 |
Rotate | .92 | .93 | .83 | .87 | .97 | .90 | .77 | .80 | .83 | .82 |
Translate | .97 | .96 | .89 | .93 | .99 | .95 | .43 | .41 | .41 | .42 |
Jumbling | .97 | .96 | .89 | .91 | .95 | .91 | .89 | .89 | .92 | .90 |
Box Jumbling | .99 | .97 | .96 | .96 | .97 | .94 | .99 | .99 | .99 | .99 |
4 Experiments
4.1 Datasets
We evaluate our method across five datasets: Kinetics-400 [25] and UCF-101 [46] for video tasks, and ImageNet-1K [8], ImageNet-Tiny-200 [29], and HMDB-51 for image tasks.
The Kinetics-400 dataset includes 240k training videos and 20k validation videos, each clip lasting 10 seconds. The dataset focuses on human-centered actions, covering a wide range of interactions. The UCF-101 dataset, while smaller with 13.3k action videos across 101 categories, offers high diversity in actions and significant variations in camera motion, object appearance, pose, background, and lighting conditions. For image-based experiments, we use the standard ImageNet-1K dataset. Additionally, we include ImageNet-Tiny-200 and the smaller video variant HMDB-51 to demonstrate our method’s effectiveness on smaller datasets, where Vision Transformers (ViTs) often struggle.
For noisy evaluation on video datasets, we assess our models under real-world perturbations following [42], along with additional noise types and intensity configurations. Video noise types include four categories: Noise (Gaussian, shot, impulse, speckle), Blur (zoom, motion, defocus), Digital (JPEG and MPEG compression artifacts), Temporal (jumble, box jumble), and Camera Motion (static rotation, dynamic rotation, translation). Each noise type is tested across five severity levels, resulting in a total of 70 unique noise configurations. In the image experiments, we evaluate model robustness using the ImageNet-P and ImageNet-C benchmarks [23]. ImageNet-C includes 15 types of corruptions, each with five severity levels, totaling 75 distinct corruptions. ImageNet-P adds 10 perturbation types with 30 intensity variations, leading to 300 different noise scenarios. All pre-training, fine-tuning, and validation are conducted on clean datasets, without noise augmentation.
4.2 Training
We use ViT-Base (ViT-B/16, input size 224) [9] as the backbone for all training and ablation studies, covering both image and video datasets. ViT-B, while over three times larger than ResNet-50 [21], is lightweight compared to other transformer architectures, making it an efficient yet powerful choice for our experiments. For image tasks, pretraining is conducted on ImageNet using a masking ratio of 0.75 and norm-pix-loss as the reconstruction objective, following the MAE framework [20]. Training spans 400 epochs with a batch size of 64, distributed across 20 nodes equipped with P100 GPUs (2 GPUs and 28 cores per node). Fine-tuning is performed for 100 additional epochs in full float-32 precision without autocast.
For video datasets, we apply the same foundational framework with adaptations tailored to temporal data. Pretraining on UCF-101 and HMDB-51 involves 800 epochs using 16-frame sequences, a batch size of 8, a tube masking ratio of 0.9, a sampling rate of 4, and a decoder depth of 4. The same node configuration is used as in the image tasks. Fine-tuning for these datasets is conducted over 100 epochs. For larger-scale datasets like Kinetics-400P, training is distributed across 12 nodes with V100 GPUs (2 GPUs and 40 cores per node) for 400 epochs, maintaining a tube masking ratio of 0.9, sampling rate of 4, and decoder depth of 4 but with an increased batch size of 16. We use DeepSpeed during video fine-tuning to manage larger batch sizes efficiently.
Our evaluation focuses on the model’s robustness to real-world noise, including shot noise, rain noise, Gaussian noise, packet loss, speckle noise, and tampering noise, as identified in prior benchmarks [37, 57]. Unlike traditional methods, we do not introduce noise augmentations during training or validation, enabling us to assess the model’s ability to generalize to unpredictable noise patterns at varying intensities. This approach mirrors real-world conditions, where noise is often diverse and unstructured. For instance, practical noise types like rain, critical for autonomous driving, and packet loss, which may affecting video-based applications. Additionally, testing on image datasets shows that our DWT-based approach is not only robust but also achieves a 2.5% accuracy improvement on clean datasets.
4.3 Evaluation
For evaluating noise robustness on Imagenet-C we compute the mean corruption error(mCE) as follows:
where, is the top-1 error for corruption and severity for the model and is the same for the baseline model for which we use AlexNet. The mean of the corruption error(mCE) across corruptions of RobutFormer and other comparable models is shown in Table 1, whereas, CE for the individual corruptions is shown in Table 3. Variants of RobustFormer methods always beats MAE [20] and CNN based approaches. Similarly, for Imagenet-P we compute the flip probability of the model ‘’ on ‘’ perturbation sequences as
where is the clean image and with are the pertubed images of . We then average across all perturbations to get mean flip error(mFP). The comparison of RobustFormer variants with Image-MAE[20] is shown in Table 2 where all the RF perform better. Figure 4, shows both the average top-1 and top-5 accuracies on Imagenet-P and Imagenet-C noise benchmarks which demonstrates that our method is robust to almost all evaluated noise types for both benchmarks.
To measure video robustness we use two metrics for relative and absolute accuracy drops. After training model , we first compute the accuracy on the clean set and accuracy for perturbation and severity . The absolute robustness and relative robustness are then computed as and . The aggregated performance for all models can be thus achieved by averaging across severity levels to get and . Table 4 shows and , which are averaged and across the perturbation categories for Kinetics-400 dataset. Our RF variants even without augmented training demonstrated significant robustness compared to some of the methods that used augmentation during training. Similarly, Table 5 shows the relative robustness scores for individual perturbations averaged across 5 severity levels. RF models here demonstrated better relative robustness to recent state-of-the-art methods. Moreover, top-1 and top-5 accuracy performance on additional noise types like rain, packetloss for UCF-101 dataset are shown in Figure 5, where RF methods are significantly superior especially in noise with high severity levels.
4.4 Ablation Study
Our preferred RobustFormer variant, RF-AA, demonstrates competitive and, in many cases, superior performance in both top-1 and top-5 accuracy compared to other configurations. Variants RF-I and RF-IA represent RobustFormer with the inclusion of the IDWT step and with both IDWT and attention steps, respectively. The IDWT step, commonly employed in traditional DWT-based methods, complements the regular DWT averaging technique.
Through our experiments, we show that incorporating the proposed DWT attention mechanism significantly enhances performance, particularly for complex noise types. For example, while the average filter alone struggles with salt-and-pepper noise, the addition of the DWT attention module leads to a marked improvement in accuracy. Similarly, the DWT attention module boosts accuracy in certain noise conditions even when applied to IDWT-based configurations, as highlighted in Figure 6 (b-d). This demonstrates the critical role of our dwt based attention mechanism in improving robustness against challenging noise scenarios.
5 Discussion and Conclusion
Noise-robust models are essential for many real-world applications, where unpredictable and unique noise patterns can significantly impact performance. While some deep learning models rely on heavy augmentation and others on rule-based filtering, these approaches are limited in scalability due to the inherent randomness and diversity of noise. RobustFormer addresses this challenge by utilizing transformers’ adaptability and the noise-handling capabilities of Wavelet Transform. This approach enables efficient decomposition of spatio-temporal information, selectively focusing on low-frequency components to enhance robustness across diverse noise types without the need for exhaustive augmentation or complex filtering rules thus making few assumptions about the evaluation noises. We validated our method on both image and video benchmark datasets with benchmark noise types for both and show our model’s significance compared to prior works.